OpenStudy (clamin):

If the standard deviation of a data set was originally 5, and if each value in the data set was multiplied by 4.8, what would be the standard deviation of the resulting data? a. 5 b. 19 c. 24 d. 14

OpenStudy (ybarrap):

Using the basic properties of variance https://en.wikipedia.org/wiki/Variance#Basic_properties $$ var(ax)=a^2var(x) $$ then $$ std(ax)=\sqrt{var(ax)}=\sqrt{a^2var(x)}=a\sqrt{var(x)}=a\times std(x) $$ So we have that $$ std(4.8x)=4.8\times std(x)=4.8\times 5 $$ Make sense?

OpenStudy (clamin):

yes

OpenStudy (clamin):

its c